OpenStudy (anonymous):

Factor. 2x2 + x − 6 (Points : 4) a. (2x − 3)(x + 2) b. (2x − 3)(x − 2) c. (2x + 3)(x − 2) d. (2x − 2)(x + 3)

OpenStudy (anonymous):

@Thehulk49

OpenStudy (er.mohd.amir):

2x2+4x-3x-6 =2x(x+2)-3(x+2) =(x+2)(2x-3)

OpenStudy (anonymous):

(x+2)(2x-3)

OpenStudy (anonymous):

will you medal the other dude please? it's only fair haha

OpenStudy (anonymous):

Wait, that's not an answer choice. It's to the power of 2 not times two

OpenStudy (anonymous):

@Thehulk49 and @Er.Mohd.AMIR

OpenStudy (anonymous):

Wait. Let me recheck mt answer.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

It's A. I think

OpenStudy (anonymous):

Because that is what I got.

OpenStudy (anonymous):

okay thank you